STIMULATED RAMAN MOLECULE FORMATION IN COLD ALKALI GASES.

\underline{PAUL JULIENNE}, Atomic Physics Division, National Institute of Standards and Technology, Gaithersburg, MD 20899; KEITH BURNETT, Clarendon Laboratory, Department of Physics, University of Oxford, Parks Road, Oxford OX1 3PU, United Kingdom; YEHUDA B. BAND, Department of Chemistry, Ben-Gurion University, Beer Sheva, Israel; WILLIAM C. STWALLEY, Department of Physics, University of Connecticut, Storrs, CT 06269.

Photoassociation is a process by which two cold, trapped atoms can be optically excited during a collision to make a vibrational level of an excited electronic state of a molecular dimer. High-resolution photoassociation spectra are now wellknown for most alkali dimers. The upper level normally decays by spontaneous emission to a number of final molecular levels, bound and unbound. A laser of a second color can be used to stimulate a transition from the upper level to a particular ground state vibrational level. Such a Raman photoassociation process can be used to form translationally cold molecules in a cold atomic gas or a Bose-Einstein condensate. Coupled channels scattering calculations show that the collisional rate coefficient for coherent production of molecules in specific target states, > 10^{−10} cm^{3}/s, can greatly exceed the rate coefficient for forming products due to excited state spontaneous decay. This implies the feasibility of rapid coherent conversion of atom pairs in a condensate to molecules on a time scale short compared to the trap oscillation time. The untrapped molecules can form a coherent molecular beam. We will examine the prospects for the experimental realization of photoassociative molecule formation in cold alkali gases or condensates.